OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint (point M) and one of the endpoints (point P). P = (3, 5) and M =(-2, 0) i got (-7, -5) am i correct? the other answers are (-1, -5) and (1/2, 5/2) thanks 4 the help:)

ganeshie8 (ganeshie8):

correct !

ganeshie8 (ganeshie8):

(3+x)/2 = -2 (5+y)2 = 0

OpenStudy (anonymous):

thank you :)

ganeshie8 (ganeshie8):

np..